### v1

# Second Robin eigenvalue bounds for Schrödinger operators on Riemannian surfaces

Abstract.
Let $(\Sigma^{2},ds^{2})$ be a compact Riemannian surface, possibly with boundary, and consider Schrödinger-type operators of the form $L=\Delta+V-aK$ together with natural Robin and Steklov-type boundary conditions incorporating a boundary potential $W$ and (in the curvature-corrected setting) the geodesic curvature $\kappa_{g}$ of $\partial\Sigma$. Our main contribution is a geometric upper bound for the second Robin eigenvalue in terms of the topology of $\Sigma$ and the integrals of $V$ and $W$, obtained via a Hersch balancing argument on the capped surface. As a geometric application, we derive sharp topological restrictions for compact two-sided free boundary minimal surfaces of Morse index at most one inside geodesic balls of negatively curved pinched Cartan–Hadamard $3$-manifolds under a mild radius condition. We also prove complementary upper bounds for first eigenvalues in the closed and Robin settings, including rigidity in the curvature-corrected case, and we establish Steklov-type estimates in a coercive regime where the Dirichlet-to-Neumann operator is well defined for all boundary data.
Key words and phrases:
Schrödinger operators, eigenvalue bounds, Robin and Steklov boundary problems, Dirichlet-to-Neumann map, free boundary minimal surfaces
2020 Mathematics Subject Classification:
Primary 35P15, 35J10, 58J50; Secondary 53A10, 49Q05
1. Introduction
Let $(\Sigma^{2},ds^{2})$ be a connected compact Riemannian surface, possibly with smooth boundary $\partial\Sigma$. Given $V\in L^{\infty}(\Sigma)$ and $a\in\mathbb{R}$, we consider the Schrödinger-type operator
(1.1) $L=\Delta+V-aK,$
where $\Delta=\mathrm{div}\,\nabla\leq 0$ is the Laplace–Beltrami operator and $K$ is the Gaussian curvature of $\Sigma$. When $\partial\Sigma\neq\emptyset$ we also allow a boundary potential $W\in L^{\infty}(\partial\Sigma)$ and a boundary operator
(1.2) $Bu=\frac{\partial u}{\partial\nu}-Wu+a\,\kappa_{g}\,u,$
where $\nu$ denotes the outward unit conormal and $\kappa_{g}$ is the geodesic curvature of $\partial\Sigma$ in $\Sigma$.
Spectral estimates for operators of the form (1.1) (and of the associated boundary problems below) are closely linked to geometry. A fundamental example is the Jacobi operator governing the second variation of area for minimal and constant mean curvature surfaces, which can be written in Schrödinger form. This connection has long been exploited to derive quantitative constraints on stability, index, and topology in surface theory; see for instance [36, 18, 12, 11, 34, 7, 6, 11, 15, 16] and the references therein.
For Schrödinger-type operators on closed surfaces, there is a large literature on upper bounds for eigenvalues in terms of conformal/topological data, going back to classical ideas of Hersch and the subsequent development of conformal volume and branched-cover methods; see e.g. [23, 38, 13, 33]. In the presence of boundary, Robin and Steklov spectra encode both interior and boundary geometry, and sharp inequalities often require combining conformal test functions with boundary trace estimates; see for instance [14, 19, 20, 10] and also [26] for related developments.
Our first main theorem is an upper bound for the second eigenvalue of the Robin problem for the general Schrödinger operator $L_{0}=\Delta+V$ (i.e. $a=0$):
(1.3) $\left\{\begin{aligned} \Delta u+Vu+\mu u&=0&&\text{in }\Sigma,\\ \partial_{\nu}u&=Wu&&\text{on }\partial\Sigma.\end{aligned}\right$.
We index Robin eigenvalues as $\mu_{0}<\mu_{1}\leq\mu_{2}\leq\cdots$, so that $\mu_{0}$ denotes the first eigenvalue.
Theorem 1.1.
Let $(\Sigma^{2},ds^{2})$ be a compact oriented surface with boundary, of genus $\gamma$. For (1.3) with $V\in L^{\infty}(\Sigma)$ and $W\in L^{\infty}(\partial\Sigma)$,
(1.4) $\mu_{1}\,|\Sigma|\;<\;8\pi\Big\lfloor\frac{\gamma+3}{2}\Big\rfloor\;-\;\int_{\Sigma}V\;-\;\int_{\partial\Sigma}W.$
The estimate (1.4) can be viewed as a boundary counterpart of the conformal topological eigenvalue bounds of El Soufi–Ilias [13] on closed manifolds (and related works), now adapted to the Robin problem on surfaces with boundary. The proof combines three ingredients: (i) we cap $\partial\Sigma$ by discs to obtain a closed surface $\widehat{\Sigma}$; (ii) we use a conformal branched cover $\widehat{\Sigma}\to\mathbb{S}^{2}$ of controlled degree (cf. [35, 9]); and (iii) we apply a Hersch balancing argument [23] to build conformal test functions, together with the min–max characterization of $\mu_{1}$.
Our second main theorem gives topological restrictions for compact two-sided free boundary minimal surfaces of low Morse index inside geodesic balls of negatively curved pinched Cartan–Hadamard $3$-manifolds. This establishes a hyperbolic analogue of the index–one topological bound proved by Chen–Fraser–Pang in ambients with $\operatorname{Ric}\geq 0$ and weakly convex boundary [9]. Such restrictions are part of a broader circle of ideas relating index bounds, curvature pinching, and topology for minimal/CMC surfaces; see for example [35, 30, 2, 8, 28] and the references therein.
Let $M^{3}$ be Cartan–Hadamard with pinched sectional curvature
$\kappa_{1}\leq\operatorname{Sec}_{M}\leq-\kappa_{2}<0,\qquad 0<\kappa_{2}\leq\kappa_{1}.$
Let $B_{o}(R)\subset M^{3}$ be a geodesic ball. If $\Sigma^{2}\subset B_{o}(R)$ is a compact two-sided free boundary minimal surface, then its Jacobi operator is $J=\Delta+V$ with $V=\operatorname{Ric}_{M}(N,N)+|A|^{2}$ and Robin boundary term $W=\mathrm{I}\!\mathrm{I}^{\partial B_{o}(R)}(N,N)$. Since $\operatorname{Ind}(\Sigma)\leq 1$ implies $\mu_{1}(J)\geq 0$, Theorem 1.1 yields restrictions once $V$ and $W$ are controlled by curvature comparison.
Theorem 1.2.
Let $M^{3}$ be Cartan–Hadamard manifold with $\kappa_{1}\leq\operatorname{Sec}_{M}\leq-\kappa_{2}<0$ and fix $o\in M$, $R>0$. Assume
(1.5) $4\kappa_{1}\tanh\!\Big(\frac{\sqrt{\kappa_{2}}\,R}{2}\Big)\ \leq\ 3\kappa_{2}\coth(\sqrt{\kappa_{2}}\,R).$
If $\Sigma^{2}\subset B_{o}(R)$ is a compact two-sided free boundary minimal surface with $\operatorname{Ind}(\Sigma)\leq 1$, then the only possibilities are
$\gamma\in\{0,1\}\ \text{ with }\ r\leq 3,\qquad\text{or}\qquad\gamma\in\{2,3\}\ \text{ with }\ r=1,$
where $\gamma$ is the genus of $\Sigma$ and $r$ is the number of boundary components.
Remark 1.3.
The left-hand side of (1.5) increases with $R$ and the right-hand side decreases with $R$, so (1.5) is a genuine geometric restriction linking $R$ and the pinching ratio $\kappa_{1}/\kappa_{2}$. We discuss its role and provide a qualitative interpretation in Section 4.
We also establish upper bounds for first eigenvalues in the closed and Robin settings, including rigidity in the curvature-corrected case (compare with intrinsic counterparts of results in immersed surface theory such as [1, 4]). Finally, we obtain Steklov-type estimates in a coercive regime where the Dirichlet problem is well posed for all boundary data and the Dirichlet-to-Neumann operator is defined on the full trace space; cf. related spectral-geometric frameworks in [14, 37, 20, 10].
Organization. Section 2 contains the closed-surface first-eigenvalue bound and a stability corollary. Section 3 proves the first Robin bound and then establishes Theorem 1.1. Section 4 proves Theorem 1.2. Section 5 develops the Jacobi–Steklov framework and proves Steklov-type eigenvalue bounds.
2. Closed surfaces: a first-eigenvalue bound
Let $\Sigma$ be a closed oriented surface and let $\lambda_{0}(L)$ denote the first eigenvalue of $L$ in (1.1). It is classical that $\lambda_{0}(L)$ is simple and admits a positive eigenfunction.
Theorem 2.1.
Let $(\Sigma^{2},ds^{2})$ be a closed oriented surface, and let $L=\Delta+V-aK$ with $V\in L^{\infty}(\Sigma)$ and $a\in\mathbb{R}$. Then
$\lambda_{0}(L)\,|\Sigma|\;\leq\;-(\operatorname*{ess\,inf}_{\Sigma}V)\,|\Sigma|+2\pi a\,\chi(\Sigma),$
where $|\Sigma|$ is the area and $\chi(\Sigma)=2-2\gamma$ is the Euler characteristic. Moreover, if $a\neq 0$ and equality holds, then $V$ and $K$ are constant and
$K=\frac{\lambda_{0}(L)+V}{a}.$
Proof.
Let $u>0$ solve $Lu+\lambda_{0}(L)u=0$. By elliptic regularity $u\in C^{\infty}(\Sigma)$, so $\ln u\in C^{\infty}(\Sigma)$. As in Perdomo [31] (see also [1]),
$\Delta\ln u=\frac{\Delta u}{u}-\frac{|\nabla u|^{2}}{u^{2}}=-\lambda_{0}(L)-V+aK-|\nabla\ln u|^{2}.$
Integrating over $\Sigma$ and using $\int_{\Sigma}\Delta\ln u=0$, we get
(2.1) $\lambda_{0}(L)|\Sigma|=-\int_{\Sigma}V+a\int_{\Sigma}K-\int_{\Sigma}|\nabla\ln u|^{2}.$
Since $V\geq\operatorname*{ess\,inf}_{\Sigma}V$ a.e. and $\int_{\Sigma}K=2\pi\chi(\Sigma)$ by Gauss–Bonnet, we obtain the desired inequality.
If equality holds and $a\neq 0$, then (2.1) forces $\nabla u\equiv 0$, so $u$ is constant, hence $V\equiv\inf_{\Sigma}V$ and $K$ is constant, yielding the formula. ∎
We say that the Schrödinger operator $L$ is stable on a closed surface if $L\geq 0$ in the $L^{2}$ sense, or equivalently $\lambda_{0}(L)\geq 0$. We have the following consequence.
Corollary 2.2.
Let $\Sigma$ be a closed oriented surface. If $L=\Delta+V-aK$ is stable on $\Sigma$, then:
  1. (1)
    for $a=0$, $\int_{\Sigma}V\leq 0$;
  2. (2)
    for $a>0$ and $V\geq 0$ a.e. on $\Sigma$, $\Sigma$ is either a sphere or a torus; in the torus case $V\equiv 0$ a.e.
Proof.
If $a=0$, from (2.1) and $\lambda_{0}(L)\geq 0$ we obtain $0\leq-\int_{\Sigma}V-\int_{\Sigma}|\nabla\ln u|^{2}$, hence $\int_{\Sigma}V\leq 0$.
If $a>0$ and $V\geq 0$, Theorem 2.1 implies $0\leq\lambda_{0}(L)|\Sigma|\leq 2\pi a\,\chi(\Sigma)$, so $\chi(\Sigma)\geq 0$, hence $\Sigma$ is a sphere or a torus. If $\chi(\Sigma)=0$ (torus), then the inequality forces $\inf_{\Sigma}V=0$. Since also $\int_{\Sigma}V\leq a\int_{\Sigma}K=0$ from (2.1), we get $V\equiv 0$ a.e. ∎
These estimates may be seen as intrinsic counterparts of results proved in the immersed setting, compare [1, Theorem 2.1] and [4, Theorem 1.1].
3. Surfaces with boundary: the Robin problem
Throughout this section, $\Sigma$ is a compact oriented surface with smooth boundary. Its Euler characteristic is
$\chi(\Sigma)=2-2\gamma-r,$
where $\gamma$ is the genus and $r$ is the number of boundary components.
3.1. The first Robin eigenvalue
Consider the Robin problem for $L=\Delta+V-aK$ with boundary operator (1.2):
(3.1) $\left\{\begin{aligned} Lu+\mu u&=0&&\text{in }\Sigma,\\ Bu&=0&&\text{on }\partial\Sigma.\end{aligned}\right$.
Its spectrum is discrete:
$\mu_{0}(L)<\mu_{1}(L)\leq\mu_{2}(L)\leq\cdots\to+\infty$
And the associated quadratic form on $H^{1}(\Sigma)$ is
(3.2) $Q[u]=\int_{\Sigma}|\nabla u|^{2}-\int_{\Sigma}(V-aK)u^{2}-\int_{\partial\Sigma}(W-a\kappa_{g})u^{2}.$
The first eigenvalue satisfies the Rayleigh characterization
(3.3) $\mu_{0}(L)=\inf_{u\in H^{1}(\Sigma)\setminus\{0\}}\frac{Q[u]}{\int_{\Sigma}u^{2}}.$
Our first estimate with boundary controls the first Robin eigenvalue $\mu_{0}(L)$ in terms of geometric and topological data of $\Sigma$, in the spirit of Theorem 2.1.
Theorem 3.1.
Let $(\Sigma^{2},ds^{2})$ be a compact oriented surface with boundary, and let $\mu_{0}(L)$ be the first eigenvalue of (3.1). Then
$\mu_{0}(L)\,|\Sigma|\;\leq\;-(\operatorname*{ess\,inf}_{\Sigma}V)\,|\Sigma|-(\operatorname*{ess\,inf}_{\partial\Sigma}W)\,|\partial\Sigma|+2\pi a\,\chi(\Sigma).$
Moreover, if $a\neq 0$ and equality holds, then $V$ and $W$ are constant a.e. and
$K=\frac{\mu_{0}(L)+V}{a},\qquad\kappa_{g}=\frac{W}{a}$
are constant.
Proof.
Test (3.3) with $u\equiv 1$. Since $V\geq\operatorname*{ess\,inf}_{\Sigma}V$ a.e. and $W\geq\operatorname*{ess\,inf}_{\partial\Sigma}W$ a.e. on $\partial\Sigma$, and by Gauss–Bonnet,
$\int_{\Sigma}K+\int_{\partial\Sigma}\kappa_{g}=2\pi\chi(\Sigma),$
we obtain the stated inequality.
If equality holds, then $u\equiv 1$ is a first eigenfunction (cf. [5, Theorem 2.4]), hence $L1+\mu_{0}(L)=0$ and $B1=0$, which imply $V$ and $W$ are constant; if $a\neq 0$ this forces $K$ and $\kappa_{g}$ to be constant as well. ∎
Corollary 3.2.
Assume $\mu_{0}(L)\geq 0$ in (3.1). If $a>0$, $V\geq 0$ on $\Sigma$, and $W\geq 0$ on $\partial\Sigma$, then $\Sigma$ is either a disk or an annulus.
Proof.
From Theorem 3.1, $0\leq\mu_{0}(L)|\Sigma|\leq 2\pi a\,\chi(\Sigma)$, hence $\chi(\Sigma)\geq 0$. Since $\chi(\Sigma)=2-2\gamma-r$ with $r\geq 1$, we must have $\gamma=0$ and $r\in\{1,2\}$. ∎
3.2. The second Robin eigenvalue: proof of Theorem 1.1
We now specialize to $a=0$, i.e. $L_{0}=\Delta+V$, and consider the Robin problem (1.3). Its quadratic form is
(3.4) $Q[u]=\int_{\Sigma}|\nabla u|^{2}-\int_{\Sigma}Vu^{2}-\int_{\partial\Sigma}Wu^{2}.$
Let $u_{0}>0$ be a first eigenfunction, and recall the min–max characterization
(3.5) $\mu_{1}=\inf\left\{\frac{Q[u]}{\int_{\Sigma}u^{2}}\,:\,u\in H^{1}(\Sigma)\setminus\{0\},\ \int_{\Sigma}u\,u_{0}=0\right\}.$
Let $\widehat{\Sigma}$ be the closed surface obtained by conformally capping each boundary component of $\Sigma$ by a disc. Then $\widehat{\Sigma}$ has the same genus $\gamma$ as $\Sigma$. A Hersch-type balancing argument on $\widehat{\Sigma}$ yields the following lemma proved by Chen, Fraser and Pang [9].
Lemma 3.3 (Chen–Fraser–Pang).
Let $u_{0}>0$ be a first eigenfunction on $\Sigma$. There exists a conformal map $\Phi:\widehat{\Sigma}\to\mathbb{S}^{2}\subset\mathbb{R}^{3}$ such that
$\int_{\Sigma}\Phi\,u_{0}=0\in\mathbb{R}^{3}\qquad\text{and}\qquad\deg(\Phi)\leq\Big\lfloor\frac{\gamma+3}{2}\Big\rfloor.$
Write $\Phi=(\phi_{1},\phi_{2},\phi_{3})$. Then $\sum_{i=1}^{3}\phi_{i}^{2}\equiv 1$ on $\Sigma$, and
(3.6) $\sum_{i=1}^{3}\int_{\Sigma}|\nabla\phi_{i}|^{2}<\sum_{i=1}^{3}\int_{\widehat{\Sigma}}|\nabla\phi_{i}|^{2}\leq 8\pi\Big\lfloor\frac{\gamma+3}{2}\Big\rfloor,$
since the Dirichlet energy of a conformal map to $\mathbb{S}^{2}$ equals $2$ times the area of the image, hence $8\pi$ times the degree, and restriction from $\widehat{\Sigma}$ to $\Sigma$ strictly decreases energy unless $\partial\Sigma=\emptyset$.
Moreover, Lemma 3.3 gives the orthogonality conditions
(3.7) $\int_{\Sigma}u_{0}\,\phi_{i}=0,\qquad i=1,2,3.$
Thus each $\phi_{i}$ is admissible in (3.5), and
(3.8) $\mu_{1}\int_{\Sigma}\phi_{i}^{2}\leq\int_{\Sigma}|\nabla\phi_{i}|^{2}-\int_{\Sigma}V\phi_{i}^{2}-\int_{\partial\Sigma}W\phi_{i}^{2}.$
Theorem 3.4.
Let $(\Sigma^{2},ds^{2})$ be a compact oriented surface with boundary of genus $\gamma$, and let $\mu_{1}$ be the second eigenvalue of (1.3). Then
$\mu_{1}\,|\Sigma|\;<\;8\pi\Big\lfloor\frac{\gamma+3}{2}\Big\rfloor\;-\;\int_{\Sigma}V\;-\;\int_{\partial\Sigma}W.$
Proof.
Sum (3.8) over $i=1,2,3$ and use $\sum_{i}\phi_{i}^{2}\equiv 1$:
$\mu_{1}|\Sigma|\leq\sum_{i=1}^{3}\int_{\Sigma}|\nabla\phi_{i}|^{2}-\int_{\Sigma}V-\int_{\partial\Sigma}W.$
Apply (3.6) to conclude the strict inequality. ∎
3.2.1. The case of planar domains
It is natural to seek upper estimates for $\mu_{1}$ that also reflect the number $r$ of boundary components. To incorporate $r$, we adapt a hemisphere mapping in the spirit of Nunes’ modified Hersch argument for free-boundary stable CMC surfaces (see [30]). By a result of Gabard [22], there exists a proper conformal branched cover
$\Phi:\Sigma\longrightarrow\mathbb{D}\cong\mathbb{S}^{2}_{+}\subset\mathbb{R}^{3},\qquad\deg\Phi\leq\gamma+r,$
and, after applying a conformal automorphism of $\mathbb{S}^{2}_{+}$ (see also [29]), we obtain the balancing conditions
(3.9) $\int_{\Sigma}u_{0}\,\phi_{i}=0,\qquad i=1,2,$
where $\Phi=(\phi_{1},\phi_{2},\phi_{3})$ with $\phi_{1}^{2}+\phi_{2}^{2}+\phi_{3}^{2}=1$ on $\Sigma$ and $\phi_{3}=0$ on $\partial\Sigma$.
Thus $\phi_{1},\phi_{2}$ are admissible in (3.5), yielding
(3.10) $\mu_{1}\int_{\Sigma}\phi_{i}^{2}\;\leq\;\int_{\Sigma}|\nabla\phi_{i}|^{2}-\int_{\Sigma}V\,\phi_{i}^{2}-\int_{\partial\Sigma}W\,\phi_{i}^{2},\qquad i=1,2.$
The Dirichlet energy is controlled by
(3.11) $\sum_{i=1}^{3}\int_{\Sigma}|\nabla\phi_{i}|^{2}\;\leq\;4\pi(\gamma+r).$
To also exploit $\phi_{3}$, which vanishes on $\partial\Sigma$, consider the Dirichlet problem
(3.12) $\left\{\begin{aligned} \Delta u+Vu+\lambda^{D}u&=0&&\text{in }\Sigma,\\ u&=0&&\text{on }\partial\Sigma,\end{aligned}\right$.
and denote its first eigenvalue by $\lambda_{0}^{D}(L_{0})$, or simply by $\lambda_{0}^{D}.$
In the case of domains in Euclidean space (thus $\gamma=0$), Filonov’s method [17] yields, in particular, the comparison
$\mu_{1}\;<\;\lambda_{0}^{D}$
for the Euclidean Laplacian with $V\equiv 0$ and $W$ with nonnegative average, which is the situation we use below.
Theorem 3.5.
Assume $\Sigma\subset\mathbb{R}^{2}$ is a bounded planar domain with $r$ boundary components, $V\equiv 0$, and $\int_{\partial\Sigma}W\geq 0$. Then
$\mu_{1}\,|\Sigma|\;<\;4\pi r\;-\;\int_{\partial\Sigma}W.$
Proof.
Using $\mu_{1}<\lambda_{0}^{D}$ and that $\phi_{3}=0$ on $\partial\Sigma$, we have
$\mu_{1}\int_{\Sigma}\phi_{3}^{2}\;<\;\int_{\Sigma}|\nabla\phi_{3}|^{2}.$
Adding this to (3.10) for $i=1,2$ and using $\phi_{1}^{2}+\phi_{2}^{2}=1$ on $\partial\Sigma$, we obtain
$\mu_{1}\int_{\Sigma}(\phi_{1}^{2}+\phi_{2}^{2}+\phi_{3}^{2})\;<\;\sum_{i=1}^{3}\!\int_{\Sigma}|\nabla\phi_{i}|^{2}-\int_{\partial\Sigma}W.$
Since $\phi_{1}^{2}+\phi_{2}^{2}+\phi_{3}^{2}\equiv 1$, the claim follows from (3.11) with $\gamma=0$. ∎
Remark 3.6.
The key extra input in the planar estimate above is the (strict) comparison $\mu_{1}<\lambda_{0}^{D}$ (in the Euclidean case, under the sign assumption on $W$), which allows us to exploit the third coordinate $\phi_{3}$ that vanishes on $\partial\Sigma$. Any extension of such a comparison to more general settings (e.g. domains in Riemannian surfaces and/or Schrödinger operators with $V\not\equiv 0$) would yield corresponding refinements of Theorem 3.5. Inequalities comparing Dirichlet and Neumann spectra of Friedlander type and their variants on Riemannian manifolds have been studied by Friedlander [21], Arendt–Mazzeo [3], Rohleder [32] and Ilias–Shouman [24], among others.
4. Application to free boundary minimal surfaces
In this section we prove Theorem 1.2. Let $M^{3}$ be complete and let $B_{o}(R)\subset M$ be a geodesic ball of radius $R$ centered at $o$. Let $\Sigma^{2}\subset B_{o}(R)$ be a compact two-sided free boundary minimal surface:
[…]
Let $N$ be a global unit normal, $A$ the shape operator of $\Sigma^{2}$, and $\mathrm{I}\!\mathrm{I}^{\partial B_{o}(R)}$ the second fundamental form of $\partial B_{o}(R)$ with respect to the outward unit normal. The index form is
$\mathcal{I}(u,u)=\int_{\Sigma}|\nabla u|^{2}-\int_{\Sigma}\big(\operatorname{Ric}_{M}(N,N)+|A|^{2}\big)u^{2}-\int_{\partial\Sigma}\mathrm{I}\!\mathrm{I}^{\partial B_{o}(R)}(N,N)\,u^{2}.$
The quadratic form $\mathcal{I}$ is precisely the Robin form associated to the Jacobi operator $J=\Delta+V$ with boundary term $W$, where
$V=\operatorname{Ric}_{M}(N,N)+|A|^{2},\qquad W=\mathrm{I}\!\mathrm{I}^{\partial B_{o}(R)}(N,N).$
Let $\{\mu_{k}(J)\}_{k\geq 0}$ be the corresponding Robin eigenvalues, ordered as $\mu_{0}(J)<\mu_{1}(J)\leq\mu_{2}(J)\leq\cdots$. We define the Morse index $\operatorname{Ind}(\Sigma)$ as the number of negative eigenvalues (counted with multiplicity). In particular,
$\operatorname{Ind}(\Sigma)\leq 1\quad\Longleftrightarrow\quad\mu_{1}(J)\geq 0.$
4.1. Curvature consequences under pinching.
Assume $M$ is a Cartan–Hadamard manifold satisfying
$\kappa_{1}\leq\operatorname{Sec}_{M}\leq-\kappa_{2}<0,\qquad 0<\kappa_{2}\leq\kappa_{1}.$
Then for any unit vector $N$,
$2\kappa_{1}\leq\operatorname{Ric}_{M}(N,N)\leq-2\kappa_{2},$
hence
$V\geq|A|^{2}-2\kappa_{1}.$
Moreover, by Hessian comparison under $\operatorname{Sec}_{M}\leq-\kappa_{2}$, the geodesic sphere $\partial B_{o}(R)$ is strictly convex and satisfies
$\mathrm{I}\!\mathrm{I}^{\partial B_{o}(R)}\ \geq\ \sqrt{\kappa_{2}}\,\coth(\sqrt{\kappa_{2}}\,R)\,g_{\partial B_{o}(R)}.$
Since $N\in T(\partial B_{o}(R))$ along $\partial\Sigma$ by the free boundary condition,
(4.1) $W=\mathrm{I}\!\mathrm{I}^{\partial B_{o}(R)}(N,N)\ \geq\ \sqrt{\kappa_{2}}\,\coth(\sqrt{\kappa_{2}}\,R)>0.$
Likewise, the geodesic curvature $\kappa_{g}$ of $\partial\Sigma$ in $\Sigma$ satisfies
(4.2) $\kappa_{g}=\mathrm{I}\!\mathrm{I}^{\partial B_{o}(R)}(T,T)\ \geq\ \sqrt{\kappa_{2}}\,\coth(\sqrt{\kappa_{2}}\,R),$
where $T$ is the unit tangent to $\partial\Sigma\subset\partial B_{o}(R)$
Finally, for a minimal surface ($H\equiv 0$), the Gauss equation gives
$K=\operatorname{Sec}_{M}(T\Sigma)-\frac{|A|^{2}}{2},$
so using $\operatorname{Sec}_{M}(T\Sigma)\geq-\kappa_{1}$ we obtain
(4.3) $|A|^{2}=2K-2\operatorname{Sec}_{M}(T\Sigma)\ \leq\ 2K+2\kappa_{1}.$
Theorem 4.1.
Let $M^{3}$ be a Cartan–Hadamard manifold whose sectional curvature satisfies
$\kappa_{1}\leq\operatorname{Sec}_{M}\leq-\kappa_{2}<0,\qquad 0<\kappa_{2}\leq\kappa_{1}.$
Fix $o\in M$ and $R>0$, and assume
(4.4) $4\kappa_{1}\tanh\!\Big(\frac{\sqrt{\kappa_{2}}\,R}{2}\Big)\ \leq\ 3\kappa_{2}\coth(\sqrt{\kappa_{2}}\,R).$
Let $\Sigma^{2}\subset B_{o}(R)$ be a compact two-sided free boundary minimal surface with $\operatorname{Ind}(\Sigma)\leq 1$. Then the only possibilities are
$\gamma\in\{0,1\}\ \text{ with }\ r\leq 3,\qquad\text{or}\qquad\gamma\in\{2,3\}\ \text{ with }\ r=1.$
Proof.
Let $\mu_{1}$ be the second Robin eigenvalue of the Jacobi operator $J=\Delta+V$ with boundary term $W$ as above. Since $\operatorname{Ind}(\Sigma)\leq 1$, we have $\mu_{1}\geq 0$. Theorem 3.4 gives
$0\leq\mu_{1}|\Sigma|<8\pi\Big\lfloor\frac{\gamma+3}{2}\Big\rfloor-\int_{\Sigma}V-\int_{\partial\Sigma}W.$
Using $V\geq|A|^{2}-2\kappa_{1}$ and (4.1),
(4.5) $0<8\pi\Big\lfloor\frac{\gamma+3}{2}\Big\rfloor-\int_{\Sigma}(|A|^{2}-2\kappa_{1})-\sqrt{\kappa_{2}}\coth(\sqrt{\kappa_{2}}R)\,|\partial\Sigma|.$
Next, from (4.3),
$\int_{\Sigma}(|A|^{2}-2\kappa_{1})=-\int_{\Sigma}|A|^{2}+2\kappa_{1}|\Sigma|\leq 2\int_{\Sigma}K+4\kappa_{1}|\Sigma|.$
Plugging into (4.5),
(4.6) $0<8\pi\Big\lfloor\frac{\gamma+3}{2}\Big\rfloor+2\int_{\Sigma}K+4\kappa_{1}|\Sigma|-\sqrt{\kappa_{2}}\coth(\sqrt{\kappa_{2}}R)\,|\partial\Sigma|.$
By Gauss–Bonnet and (4.2),
$\int_{\Sigma}K=2\pi(2-2\gamma-r)-\int_{\partial\Sigma}\kappa_{g}\leq 2\pi(2-2\gamma-r)-\sqrt{\kappa_{2}}\coth(\sqrt{\kappa_{2}}R)\,|\partial\Sigma|.$
Substituting into (4.6) yields
(4.7) $0<8\pi\Big\lfloor\frac{\gamma+3}{2}\Big\rfloor+4\pi(2-2\gamma-r)+4\kappa_{1}|\Sigma|-3\sqrt{\kappa_{2}}\coth(\sqrt{\kappa_{2}}R)\,|\partial\Sigma|.$
Finally, apply the linear isoperimetric inequality of Lee–Seo [25, Theorem 2.10(ii)] (with $n=2$, $k=\sqrt{\kappa_{2}}$, and $H=0$):
$|\Sigma|\leq\frac{1}{\sqrt{\kappa_{2}}}\tanh\!\Big(\frac{\sqrt{\kappa_{2}}R}{2}\Big)\,|\partial\Sigma|.$
Therefore,
$4\kappa_{1}|\Sigma|\leq\frac{4\kappa_{1}}{\sqrt{\kappa_{2}}}\tanh\!\Big(\frac{\sqrt{\kappa_{2}}R}{2}\Big)\,|\partial\Sigma|.$
Substitute into (4.7) and use (4.4) to drop the resulting nonpositive boundary-length term. We obtain
$2\Big\lfloor\frac{\gamma+3}{2}\Big\rfloor+2-2\gamma-r>0.$
Writing $\gamma=2m$ or $\gamma=2m+1$ shows $2\lfloor(\gamma+3)/2\rfloor+2-2\gamma=4-2m$, hence 42mr>04-2m-r>0, i.e. $r\leq 3-2m$. Thus: $m=0$ gives $\gamma\in\{0,1\}$ and $r\leq 3$, while $m=1$ gives $\gamma\in\{2,3\}$ and $r\leq 1$, which completes the proof. ∎
Remark 4.2.
Since $\tanh(\sqrt{\kappa_{2}}R/2)\nearrow 1$ and $\coth(\sqrt{\kappa_{2}}R)\searrow 1$ as $R\to\infty$, the limiting form of (4.4) would be $4\kappa_{1}\leq 3\kappa_{2}$. Under our convention $\kappa_{1}\geq\kappa_{2}$, this cannot hold; hence (4.4) is necessarily a small-radius requirement. For $R\to 0$, one has $\tanh(\sqrt{\kappa_{2}}R/2)\sim\sqrt{\kappa_{2}}R/2$ and $\coth(\sqrt{\kappa_{2}}R)\sim 1/(\sqrt{\kappa_{2}}R)$, so (4.4) holds for $R$ sufficiently small (depending on $\kappa_{1},\kappa_{2}$).
5. A Jacobi–Steklov framework
In this section we study the Steklov-type eigenvalue problem associated with the pair $(L_{0},B_{0})$:
(5.1) $\left\{\begin{array}[]{rcll}L_{0}u&=&0&\text{in }\Sigma,\\[2.84526pt] B_{0}u&=&\sigma u&\text{on }\partial\Sigma,\end{array}\right.\qquad L_{0}=\Delta+V,\quad B_{0}=\partial_{\nu}-W,$
where $\Sigma$ is a compact oriented surface with smooth boundary and $V\in L^{\infty}(\Sigma)$, $W\in L^{\infty}(\partial\Sigma)$. Related Steklov-type problems arising from free boundary minimal surface geometry are discussed, for instance, by Lima–Menezes [27].
5.1. Coercivity and the Dirichlet-to-Neumann map
Unlike the classical Steklov problem (where $V\equiv 0$ and $W\equiv 0$), the Dirichlet problem for $L_{0}$ may fail to be solvable for arbitrary boundary data. Consequently, the corresponding Dirichlet-to-Neumann map may only be defined on a proper subspace of boundary traces. This issue is analyzed in detail by Tran [37] in a slightly more general framework. To avoid such domain restrictions, we work in a coercive regime that guarantees well-posedness for all boundary data, namely positivity of the first Dirichlet eigenvalue.
Consider the Dirichlet problem
$\left\{\begin{aligned} \Delta u+Vu+\lambda^{D}u&=0&&\text{in }\Sigma,\\ u&=0&&\text{on }\partial\Sigma.\end{aligned}\right$.
Let $\lambda_{0}^{D}(L_{0})$ denote its first eigenvalue.
Assume henceforth that $\lambda_{0}^{D}(L_{0})>0$. Then the interior quadratic form
$Q^{D}[u]:=\int_{\Sigma}|\nabla u|^{2}-\int_{\Sigma}Vu^{2}$
is coercive on $H_{0}^{1}(\Sigma)$. In particular, for each boundary datum $h\in H^{1/2}(\partial\Sigma)$ there exists a unique minimizer of $Q^{D}$ among all $H^{1}$-extensions of $h$. This minimizer is precisely the weak $L_{0}$-harmonic extension of $h$, and it provides a canonical definition of the Dirichlet-to-Neumann map.
Proposition 5.1.
Assume $\lambda_{0}^{D}(L_{0})>0$. Given $h\in H^{1/2}(\partial\Sigma)$, consider the affine space
$\mathcal{A}_{h}:=\{u\in H^{1}(\Sigma):u|_{\partial\Sigma}=h\}.$
Then there exists a unique $\widehat{h}\in\mathcal{A}_{h}$ such that
$Q^{D}[\widehat{h}]=\min_{u\in\mathcal{A}_{h}}Q^{D}[u].$
Moreover, $\widehat{h}$ is the unique weak solution of $L_{0}u=0$ in $\Sigma$ with trace $h$.
Proof.
Fix $h\in H^{1/2}(\partial\Sigma)$ and choose $u_{h}\in H^{1}(\Sigma)$ with $u_{h}|_{\partial\Sigma}=h$. For $v\in H_{0}^{1}(\Sigma)$ set $u=u_{h}+v$, so $u\in\mathcal{A}_{h}$.
Since $\lambda_{0}^{D}(L_{0})>0$, the form $Q^{D}$ is coercive on $H_{0}^{1}(\Sigma)$. Hence the functional $v\mapsto Q^{D}[u_{h}+v]$ is strictly convex, coercive, and weakly lower semicontinuous on $H_{0}^{1}(\Sigma)$, and therefore it admits a unique minimizer $v_{*}\in H_{0}^{1}(\Sigma)$. Setting $\widehat{h}:=u_{h}+v_{*}$ yields the unique minimizer of $Q^{D}$ in $\mathcal{A}_{h}$.
The Euler–Lagrange equation gives, for every $\varphi\in H_{0}^{1}(\Sigma)$,
$\frac{d}{dt}\Big|_{t=0}Q^{D}[\widehat{h}+t\varphi]=0\quad\Longleftrightarrow\quad\int_{\Sigma}\langle\nabla\widehat{h},\nabla\varphi\rangle-\int_{\Sigma}V\,\widehat{h}\,\varphi=0,$
which is exactly the weak formulation of $L_{0}\widehat{h}=0$ with trace $h$. Uniqueness follows from coercivity. ∎
Under Proposition 5.1, we define the Dirichlet-to-Neumann map $\Lambda_{L_{0}}:H^{1/2}(\partial\Sigma)\to H^{-1/2}(\partial\Sigma)$ by
$\Lambda_{L_{0}}(h):=\left.\partial_{\nu}\widehat{h}\right|_{\partial\Sigma},$
and the associated boundary operator
$\mathcal{S}:=\Lambda_{L_{0}}-W.$
Then (5.1) is equivalent to $\mathcal{S}h=\sigma h$ on $\partial\Sigma$, where $h=u|_{\partial\Sigma}$. In the coercive regime $\lambda_{0}^{D}(L_{0})>0$, the operator $\mathcal{S}$ defines a self-adjoint operator on $L^{2}(\partial\Sigma)$ with discrete spectrum, and its eigenvalues coincide with the Steklov spectrum of (5.1) (cf. [37]).
5.2. Variational characterizations
Let
$Q[u]=\int_{\Sigma}|\nabla u|^{2}-\int_{\Sigma}Vu^{2}-\int_{\partial\Sigma}Wu^{2},\qquad u\in H^{1}(\Sigma).$
If $u$ is $L_{0}$-harmonic in $\Sigma$, then integration by parts yields
(5.2) $Q[u]=\int_{\partial\Sigma}u(\partial_{\nu}u-Wu)=\int_{\partial\Sigma}u\,B_{0}u.$
In particular, if $u$ solves (5.1), then $Q[u]=\sigma\int_{\partial\Sigma}u^{2}$.
Using harmonic extensions $\widehat{h}$ from Proposition 5.1, we obtain the Rayleigh–Ritz characterizations
(5.3) $\sigma_{0}=\inf_{h\in H^{1/2}(\partial\Sigma)\setminus\{0\}}\frac{Q[\widehat{h}]}{\int_{\partial\Sigma}h^{2}},$
and, if $h_{0}$ is the boundary trace of a first Steklov eigenfunction,
(5.4) $\sigma_{1}=\inf\left\{\frac{Q[\widehat{h}]}{\int_{\partial\Sigma}h^{2}}:\ h\in H^{1/2}(\partial\Sigma)\setminus\{0\},\ \int_{\partial\Sigma}h\,h_{0}=0\right\}.$
Remark 5.2.
Without assuming $\lambda_{0}^{D}(L_{0})>0$, one can still define a Dirichlet-to-Neumann map for $L_{0}$ on a suitable closed subspace of $L^{2}(\partial\Sigma)$ (and hence a corresponding Steklov spectrum), exactly as in Tran [37]. We restrict to the coercive regime to avoid introducing these domain restrictions.
5.3. An upper bound for the first Steklov eigenvalue in the curvature-corrected case
We return to the curvature-corrected pair
$L=\Delta+(V-aK),\qquad B=\partial_{\nu}-W+a\kappa_{g},$
and assume $\lambda_{0}^{D}(L)>0$.
Theorem 5.3.
Let $(\Sigma^{2},ds^{2})$ be a compact oriented surface with boundary and assume $\lambda_{0}^{D}(L)>0$. Then
$\sigma_{0}\leq-\frac{1}{|\partial\Sigma|}\int_{\Sigma}V-\frac{1}{|\partial\Sigma|}\int_{\partial\Sigma}W+\frac{2\pi a\,\chi(\Sigma)}{|\partial\Sigma|}.$
Consequently,
(5.5) $\sigma_{0}\leq-\frac{|\Sigma|}{|\partial\Sigma|}\,\operatorname*{ess\,inf}_{\Sigma}V-\operatorname*{ess\,inf}_{\partial\Sigma}W+\frac{2\pi a\,\chi(\Sigma)}{|\partial\Sigma|}.$
Moreover, if $a\neq 0$ and equality holds in (5.5), then $V$ and $W$ are constant a.e. and
$K=\frac{V}{a},\qquad\kappa_{g}=\frac{\sigma_{0}+W}{a}$
are constant.
Proof.
Let $h\equiv 1$ on $\partial\Sigma$ and let $\widehat{1}$ be its $L$-harmonic extension. By (5.3),
$\sigma_{0}\leq\frac{Q[\widehat{1}]}{\int_{\partial\Sigma}1}=\frac{Q[\widehat{1}]}{|\partial\Sigma|}.$
By Proposition 5.1 applied to $L$, the harmonic extension $\widehat{1}$ minimizes the interior energy $Q^{D}$ among all extensions with trace $1$, hence $Q^{D}[\widehat{1}]\leq Q^{D}[1]$. Since the boundary term in $Q$ is fixed by the trace, we obtain $Q[\widehat{1}]\leq Q[1]$, and therefore $\sigma_{0}\leq Q[1]/|\partial\Sigma|$. Finally, by the definition of $Q$ and Gauss–Bonnet,
$Q[1]=-\int_{\Sigma}(V-aK)-\int_{\partial\Sigma}(W-a\kappa_{g})=-\int_{\Sigma}V-\int_{\partial\Sigma}W+2\pi a\,\chi(\Sigma),$
which yields the first inequality; (5.5) follows from the $\operatorname*{ess\,inf}$ bounds.
If equality holds in (5.5), then $Q[\widehat{1}]=Q[1]$, and by uniqueness of the $L$-harmonic extension this forces $\widehat{1}\equiv 1$. Hence $L1=0$ in $\Sigma$ and $B1=\sigma_{0}$ on $\partial\Sigma$, implying $V=aK$ and $W+a\kappa_{g}=\sigma_{0}$ a.e. If $a\neq 0$, this yields the stated rigidity. ∎
5.4. An upper bound for the second Steklov eigenvalue
We state an estimate in the coercive regime $\lambda_{0}^{D}(L_{0})>0$, obtained by combining the Fraser–Schoen [19] conformal mapping method with the minimizing property of harmonic extensions. We emphasize that the key conformal input is the same hemisphere mapping used earlier in Subsubsection 3.2.1 (cf. (3.11)); the difference is that here the min–max constraint is expressed in terms of boundary traces, so the balancing condition is imposed on $\partial\Sigma$.
Theorem 5.4.
Let $(\Sigma^{2},ds^{2})$ be a compact oriented surface with boundary of genus $\gamma$ and $r$ boundary components. Assume $\lambda_{0}^{D}(L_{0})>0$ for $L_{0}=\Delta+V$, and let $\sigma_{1}=\sigma_{1}(L_{0},B_{0})$ be the second Steklov eigenvalue of (5.1) with $B_{0}=\partial_{\nu}-W$. Then
$\sigma_{1}\,|\partial\Sigma|<4\pi(\gamma+r)-\int_{\Sigma}V-\int_{\partial\Sigma}W.$
Proof.
Let $u_{0}$ be a first Steklov eigenfunction and set $h_{0}=u_{0}|_{\partial\Sigma}$. By Gabard’s theorem [22], there exists a proper conformal branched cover $\Phi=(\phi_{1},\phi_{2},\phi_{3}):\Sigma\to\mathbb{S}^{2}_{+}\subset\mathbb{R}^{3}$ of degree at most $\gamma+r$ mapping $\partial\Sigma$ into the equator. In particular,
$\sum_{i=1}^{3}\int_{\Sigma}|\nabla\phi_{i}|^{2}\leq 4\pi(\gamma+r),\qquad\phi_{3}|_{\partial\Sigma}=0,\qquad\phi_{1}^{2}+\phi_{2}^{2}=1\ \text{on }\partial\Sigma.$
Postcomposing with a conformal automorphism of $\mathbb{S}^{2}_{+}$, we may assume
$\int_{\partial\Sigma}\phi_{1}h_{0}=\int_{\partial\Sigma}\phi_{2}h_{0}=0,$
by the Hersch balancing argument as in Fraser–Schoen [19]. Let $\widehat{\phi}_{i}$ be the $L_{0}$-harmonic extension of $\phi_{i}|_{\partial\Sigma}$. Then $\widehat{\phi}_{i}$ is admissible in (5.4) and, by the minimizing property of harmonic extensions,
$Q[\widehat{\phi}_{i}]\leq Q[\phi_{i}],\qquad i=1,2.$
Therefore,
$\sigma_{1}\int_{\partial\Sigma}\phi_{i}^{2}\leq Q[\phi_{i}],\qquad i=1,2.$
Summing and using $\phi_{1}^{2}+\phi_{2}^{2}=1$ on $\partial\Sigma$ yields
$\sigma_{1}|\partial\Sigma|\leq Q[\phi_{1}]+Q[\phi_{2}].$
Since $\phi_{3}|_{\partial\Sigma}=0$, the Dirichlet variational principle gives
$\int_{\Sigma}V\phi_{3}^{2}\leq\int_{\Sigma}|\nabla\phi_{3}|^{2}-\lambda_{0}^{D}(L_{0})\int_{\Sigma}\phi_{3}^{2}.$
Using $\phi_{1}^{2}+\phi_{2}^{2}=1-\phi_{3}^{2}$ on $\Sigma$, one finds
$Q[\phi_{1}]+Q[\phi_{2}]\leq\sum_{i=1}^{3}\int_{\Sigma}|\nabla\phi_{i}|^{2}-\int_{\Sigma}V-\int_{\partial\Sigma}W-\lambda_{0}^{D}(L_{0})\int_{\Sigma}\phi_{3}^{2}.$
Combining with the energy bound gives
$\sigma_{1}|\partial\Sigma|+\lambda_{0}^{D}(L_{0})\int_{\Sigma}\phi_{3}^{2}\leq 4\pi(\gamma+r)-\int_{\Sigma}V-\int_{\partial\Sigma}W.$
Since $\lambda_{0}^{D}(L_{0})>0$ and $\Phi$ is nonconstant, we have $\phi_{3}\not\equiv 0$, hence the extra term is strictly positive, yielding the strict inequality. ∎
In the proof above we used a branched conformal cover $\Phi=(\phi_{1},\phi_{2},\phi_{3}):\Sigma\to\mathbb{S}^{2}_{+}$ in order to exploit the boundary identity $\phi_{1}^{2}+\phi_{2}^{2}=1$ on $\partial\Sigma$, at the expense of the energy bound $\sum_{i=1}^{3}\int_{\Sigma}|\nabla\phi_{i}|^{2}\leq 4\pi(\gamma+r)$. If one assumes in addition that $V\leq 0$ on $\Sigma$, one can instead work directly with Gabard’s branched cover $\varPhi=(\varphi^{1},\varphi^{2}):\Sigma\to\mathbb{D}$ of degree at most $\gamma+r$. In this case $(\varphi^{1})^{2}+(\varphi^{2})^{2}=1$ on $\partial\Sigma$ and $(\varphi^{1})^{2}+(\varphi^{2})^{2}\leq 1$ on $\Sigma$, and the nonpositivity of $V$ allows the interior deficit $1-(\varphi^{1})^{2}-(\varphi^{2})^{2}$ to be handled without introducing a third coordinate. This yields the sharper constant $2\pi(\gamma+r)$.
Corollary 5.5.
Let $(\Sigma^{2},ds^{2})$ be a compact oriented surface with boundary of genus $\gamma$ and $r$ boundary components. Assume $\lambda_{0}^{D}(L_{0})>0$ for $L_{0}=\Delta+V$, and suppose moreover that $V\leq 0$ on $\Sigma$. Let $\sigma_{1}=\sigma_{1}(L_{0},B_{0})$ be the second Steklov eigenvalue of (5.1) with $B_{0}=\partial_{\nu}-W$. Then
$\sigma_{1}\,|\partial\Sigma|\leq 2\pi(\gamma+r)-\int_{\Sigma}V-\int_{\partial\Sigma}W.$
Acknowledgments
We thank Ivaldo Nunes and João Henrique Andrade for their interest in this work and for valuable comments and suggestions.
This work was carried out within the CAPES–Math AmSud project New Trends in Geometric Analysis (CAPES, Grant 88887.985521/2024-00).
R. A. was partially supported by the Brazilian National Council for Scientific and Technological Development (CNPq, Grant 151804/2024-9). M. C. was partially supported by CNPq (Grants 405468/2021-0 and 311136/2023-0). V. S. was partially supported by FAPEAL (Grant E:60030.0000000971/2024).